ETHYL ALCOHOL A CASE STUDY IN ASYMMETRIC-ASYMMETRIC INTERNAL ROTATION

\underline{J. C. PEARSON}, Jet Propulsion Laboratory, California Institute of Technology, Mail Stop 301-429, 4800 Oak Grove Dr., Pasadena, CA 91109.

Ethyl Alcohol offers an outstanding opportunity to study the effects of Asymmetric-Top on an Asymmetric-Frame internal rotation. The ground state spectrum of ethyl alcohol is comprised of three torsional substates; the trans or e_{0}, the symmetric gauche (gauche+ or e_{1}) and the asymmetric gauche (gauche- or o_{1}). The interactions and transitions are governed by the C_{S} group, with all symmetries being either odd or even. Ethyl alcohol is a prolate asymmetric top and the trans substate has slightly larger rotational constants than the two gauche substates. The trans substate is 1185.4 GHz below the gauche+ state. As a result of the structure and torsional energy, the energy level structure is rich in strongly avoided crossings and substantial mixing, often of all three substates. The spectrum has been assigned to J = 70 and K_{a} = 24, including multiple avoided crossings between trans and gauche states with K_{a} differences of less than 3. The ethyl alcohol energy level structure and spectral features unique to the asymmetric internal rotation problem will be described.